A train slowing to a stop has an average acceleration of -3.00 m/ s 2 . If its initial velocity is 30.0 m/s, how far does it travel in 4.00 seconds?

A.) 144m
B.) 140m
C.) 96m
D.) 96.0m
Answer #1

шля using the joumula s = ut t & at² S = 30x4 + 1 X (-307 s = 120-24 is = 96m so option (C) should be right.